I need help solving using substitution method
3x + 24y = 24
27x - 15y = -15y

Answer by stanbon(75887)

solving using substitution method
3x + 24y = 24
27x - 15y = -15
-----------------------
Simplify each equation:
x + 8y = 8
9x - 5y = -5
----
Modify:
x = 8-8y
------
Substitute:
9(8-8y)-5y = -5
72-72y-5y = -5
77y = 77
y = 1
-----
Since x = 8-8y, x = 0
============================
Solution: (0,1)
======================
Cheers,
Stan H.
===========